# A Centralized Planning and Distributed Execution Method for Shape Filling with Homogeneous Mobile Robots

### v1

@@ C.7.2 Localization for Recycled robots
Theorem 37.
For epoch $N+1\leq T\leq 2N$, robot $u_{N+k}$ executing the edge-following movement stops at $t_{r}[k]$.
Proof.
By lemma 34, lemma 35 and lemma 36, the theorem follows. ∎
@@ C.8 Conclusion
    For each prefix subsequence $(a_{1},b_{1},1),\ldots,(a_{k},b_{k},k)$, $1\leq k\leq 2N$, let $A_{k}=\{a_{1},\ldots,a_{k-1}\}$, $B_{k}=\{b_{1},\ldots,b_{k-1}\}$, and $\overline{B_{k}}=B_{k}$ since the add-subtract algorithm enforce at most one moving robot at any time. Then the starting position of the $k$th robot $a_{k}\in I\oplus\overline{B_{k}}\circleddash A_{k}$ and the robot can only relay on robots centered at $I\oplus\overline{B_{k}}\circleddash A_{k}\circleddash\{a_{k}\}$ to locate its position.
@@
As a result, $\{\{b_{1},\ldots,b_{2N}\}\circleddash\{a_{1},\ldots,a_{2N}\}\}\sqcap H_{1}=\{S\oplus\{b_{N+1},\ldots,b_{2N}\}\circleddash I\circleddash S\}\sqcap H_{1}=S\setminus D$. This shows theorem 38.1. For $1\leq k\leq N$, $a_{k}\in I\circleddash\{a_{1},\ldots,a_{k-1}\}\subseteq I\oplus\overline{B_{k}}\circleddash A_{k}$. For $N+1\leq k\leq 2N$, $a_{k}\in S\circleddash\{{t_{r}}_{1},\ldots,{t_{r}}_{k-N}\}\subseteq I\oplus\overline{B_{k}}\circleddash A_{k}$. Furthermore, notice that the active robots are localizable by proposition 18, proposition 35 and proposition 36. Theorem 38.2 follows.
Finally, by theorem 19 and theorem 37, the robot starts from lattice position $a_{k}$ stops at the stopping lattice position $b_{k}$ by executing the add-subtract algorithm. This completes the proof. ∎

### v3

@@ C.8 Conclusion
    For each prefix subsequence $(a_{1},b_{1},1),\ldots,(a_{k},b_{k},k)$, $1\leq k\leq 2N$, let $A_{k}=\{a_{1},\ldots,a_{k-1}\}$, $B_{k}=\{b_{1},\ldots,b_{k-1}\}$, and $\overline{B_{k}}=B_{k}$ since the add-subtract algorithm enforce at most one moving robot at any time. Then the starting position of the $k$th robot $a_{k}\in I\oplus\overline{B_{k}}\circleddash A_{k}$ and the robot can only relay on robots centered at $I\oplus\overline{B_{k}}\circleddash A_{k}\circleddash\{a_{k}\}$ to locate its position before stopping at $b_{k}$.
@@
As a result, $\{\{b_{1},\ldots,b_{2N}\}\circleddash\{a_{1},\ldots,a_{2N}\}\}\sqcap H_{1}=\{S\oplus\{b_{N+1},\ldots,b_{2N}\}\circleddash I\circleddash S\}\sqcap H_{1}=S\setminus D$. This shows theorem 36.1. For $1\leq k\leq N$, $a_{k}\in I\circleddash\{a_{1},\ldots,a_{k-1}\}\subseteq I\oplus\overline{B_{k}}\circleddash A_{k}$. For $N+1\leq k\leq 2N$, $a_{k}\in S\circleddash\{{t_{r}}_{1},\ldots,{t_{r}}_{k-N}\}\subseteq I\oplus\overline{B_{k}}\circleddash A_{k}$. Furthermore, by lemma 18, lemma 34 and lemma 35, the active robots are localizable whenever necessary. By lemma 15 and lemma 33, the path for each moving robot is unobstructed. Theorem 36.2 follows.